OpenStudy (anonymous):

When a 1.0 kg object rest on your foot for 0.10 s, the impulse exerted is -0.98 Ns. It's t/f. The formula for impulse is (Please correct me if I'm wrong) force=(mass*velocity)/time So wouldn't it be force=(1.0*9.8)/0.10 = 9.8/0.10 = 98. Right ? How do you know if it's supposed to be negative ? Or is it negative because it is falling ? I don't know.. Thanks for helping though !

OpenStudy (noelgreco):

\[Impulse = F \Delta t = \Delta p\] Imagine kicking a football. The ball does not really "rest" on your foot, but is in contact with it for a certain time. The impulse of the ball on your foot is negative; the impulse of your foot on the ball is positive. "So wouldn't it be force=(1.0*9.8)/0.10 = 9.8/0.10 = 98." No, it would be 1.0 * 9.8 * 0.10 = .98. Add the negative depending on the direction of the impulse.

OpenStudy (anonymous):

Ohh. Okay. I understand. Thanks, the football made it easy to see what your saying. Thanks !